OpenStudy (anonymous):

approximate (-1,3) PLEASE HELP!!!!!

OpenStudy (richyw):

there is nothing to approximate. assuming those are a coordinate pair? do you want to approximate its distance from the origin?

OpenStudy (anonymous):

the x-axis on the coordinate plane is what I need to find out

OpenStudy (anonymous):

need to round off to the nearest hundredth

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

the assignment is about the game angry birds. need to approximate the x-axis on a coordinate plane for the equation \[y= -x^2 - 2x + 6\]

OpenStudy (richyw):

I really don't understand what you mean by approximate the x axis on a coordinate plane? Is that exactly what the question asked?

OpenStudy (richyw):

do you want the x intercepts?

OpenStudy (anonymous):

the question says: suppose that the pig is sitting on the x axis of a coordinate plane. On a coordinate plane, where is the pig on the x axis, assuming that the bird is flying left to right? (Round to the nearest hundredth)

OpenStudy (richyw):

ok this is not enough information. If this is all you have on your assignment, then you cannot answer the question. However I suspect that you are leaving information out still. If the assignment is telling you that the angry bird is launched from one x intercept of the parabola, and hits the pig on the other x intercept, then it is easy to find the pigs coordinates, but this has nothing to do with the point (-1,3). I have a strong suspicion that all you must to is find the positive x intercept of the function you have provided.

OpenStudy (anonymous):

ok, how do I go about finding that ?

OpenStudy (anonymous):

I know that it starts at the beginning of the arch and ends at the end, but not sure how to find that in hundredths

OpenStudy (richyw):

the x intercepts of the function are at the points where \(y=0\) so set that equation to zero and solve. you will find that this function is not easily factored so you will need to plug it into the quadratic equation to find its roots. The larger of the two roots will be your exact answer. If you want to approximate that root to the nearest hundreth this question becomes a lot harder, if you can use a calculator then it is simple. The quadratic equation is \[\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (richyw):

that equation is for the polynomial \[ax^2+bx+c\]

OpenStudy (anonymous):

this is the whole worksheet that I am working with. maybe this will help

OpenStudy (anonymous):

its a quadratic equation \[x = b/2a\]

OpenStudy (richyw):

ok that document does not give enough information for a lot of the questions since the coordinate of the pigs is not given. However for the question about the pig on the x axis, if you assume that the pig is in the flight path of the bird, then the procedure i explaned above will get you the correct answer

OpenStudy (anonymous):

ok thank you for all of your help. wish me luck cause I'm gonna need it! Thanks again

OpenStudy (richyw):

no worries. Once you have done the work, check and make sure your answer is \(\sqrt{7}-1\). Which you can hopefully just use a calculator to approximate.